Write a Python program (using python 3.7.2) that lets the user play the game of Rock, Paper, Scissors against the computer.

The program should work as follows.

1. When the program begins, a random number in the range of 1 through 3 is generated. If the

number is 1, then the computer has chosen rock. If the number is 2, then the computer has

chosen paper. If the number is 3, then the computer has chosen scissors. (Don’t display the

computer’s choice yet.)

2. The user enters his

or her choice of “rock”, “paper”, or “scissors” at the keyboard. (You can use

a menu if you prefer.)

3. The computer’s choice is displayed.

4. A winner is selected according to the following rules:

a. The rock smashes the scissors.

b. Scissors cuts paper.

c. Paper wraps rock.

d. If both players make the same choice it is a tie.

Print who the winner is.

Part B:

Once you have completed Part A continue playing, keeping track of how many wins, losses and ties. Add

an item to the menu that asks if the user wants to Quit. Keep playing until the user selects Quit. Print out the totals.

USE COMMENTS.

Notes: To generate a random integer 1 ,2 or 3:

Import random

num=random.randint(1,3)
Answer #1

Code:-

import random #importing libraries
win=0 #initially declare win with 0
lose=0 #initially declare lose with 0
tie=0 #initially declare lose with 0
while(True): #infinite loop
num=random.randint(1,3) #chose random number between 1 and 3
user=int(input("Enter Your Choice 1)rock 2)paper 3)scissors 4)quit: ")) #take input from user menu driven
if user==1: #if user selects rock
if num==1: #if computer chooses rock
print("Computer chooses rock")
print("Both chooses rock")
tie=tie+1
if num==2: #if computer chooses paper
print("Computer chooses paper")
print("paper wraps rock user loses")
lose=lose+1
if num==3: #if computer chooses scissors
print("Computer chooses scissors")
print("rock smashes scissors user wins")
win=win+1
  
elif user==2: #if user selects paper
if num==2: #if computer chooses paper
print("Computer chooses paper")
print("Both chooses paper")
tie=tie+1
if num==3: #if computer chooses scissors
print("Computer chooses scissors")
print("Scissors cuts paper user loses")
lose=lose+1
if num==1: #if computer chooses rock
print("Computer chooses rock")
print("paper wraps rock user wins")
win=win+1
elif user==3: #if user selects scissors
if num==3: #if computer chooses scissors
print("Computer chooses scissors")
print("Both chooses scissors")
tie=tie+1
if num==1: #if computer chooses rock
print("Computer chooses rock")
print("rock smashes scissors user loses")
lose=lose+1
if num==2: #if computer chooses paper
print("Computer chooses paper")
print("Scissors cuts paper user wins")
win=win+1
elif user==4: #if user want to quit
break
else: #if choice is invalid
print("Invalid choice")
print("Total Wins",win) #printing total wins
print("Total loses",lose) #printing total loses
print("Total ties",tie) #printing total ties

Code Screen Shot:-

Output:-

#Any Doubts please comment